40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1470     Entry time: Fri Apr 10 18:11:18 2009
Author: Jenne 
Type: Update 
Category: PSL 
Subject: ISS has a bad cable? 

[Rob, Jenne]

I noticed that the ISS Mean Value and CS Saturation were both RED and unhappy. (The alarms were going off, and they were both red on the MEDM screen).  None of the MEDM settings seemed off kilter, so we went out to take a look at the PSL table. 

Rob checked that light is indeed going to both of the ISS photodiodes (Morag and Siobhan).  Next we checked that all the cables were good, and that the power to the ISS box was plugged in. In this process, Rob wiggled all the cables to check that they were plugged in.  Just after doing this, the Mean Value and CS Sat were happy again.  Rob thinks the current shunt connection might be bad, but we don't really know which one it was since all of the cables were jiggled between our checking the screens. 

Right now, everything is happy again, but as with all bad-cabling-problems, we'll probably see this one again.

 

 

I don't know why in particular the connection decided to spaz out this afternoon...I don't think anyone opened the PSL table before Rob and I went to investigate.  I was working on the PMC servo (checking the LO levels...to be posted in a couple minutes), but didn't have anything to do with the ISS. After I was done, I put everything back, and locked the PMC and the MC, and everything was good, until some time later when the ISS started flipping out.

ELOG V3.1.3-